A mixture of gas is two moles of hydrogen and four moles of oxygen was a total pressure of 3atm. a. calculate the mole fraction of the hydrogen in the mixture. b. calculate the partial pressure in atm of hydrogen and oxygen gas in the mixture.

XH2=

PH2=

PO2=
Answer #1

MH₂ = 2 nog = 4. mole traction of H₂ = 2/2 = 2 = 0.3333 mole traction of Op = 4 = 0.6666 244 4+2 Partial pressure of Og = mol